OpenStudy (anonymous):

multiply (6+4i)(9+i)=

OpenStudy (zehanz):

Have you heard of FOIL? The only extra thing you must remember is: wherever there's an i², replace it with -1.

OpenStudy (zehanz):

First, Outer, Inner, Last

OpenStudy (anonymous):

no

OpenStudy (anonymous):

If you do FOIL method, you end up with: 54 + 6i+ 36i+ 4i^2 Combine like terms, you get 42i+54+4i^2 When there is an i^2, it = -1. So if you do 4(-1), you get -4 54-4+42i 50+42i

OpenStudy (zehanz):

You must multiply every number in the left pair of brackets with each one in the right pair: First: 6*9 Outer:6*i Inner:4i*9 Last: 4i*i Add everything up where possible, and remember: i²=-1

OpenStudy (anonymous):

problem hav been solved?

OpenStudy (anonymous):

im having trouble getting the answer

OpenStudy (anonymous):

do u study complex number?

OpenStudy (zehanz):

As @zaynahf calculated: First: 54 Outer: 6i Inner: 36i Last: 4i² = 4*-1 = -4 Now complex numbers can be written as a+bi, a is the real part and b is the imaginary part. Here we have 54 and -4, they add up to 50. 6i+36i = 42i. End answer: 50 + -4i = 50-4i

OpenStudy (anonymous):

no

OpenStudy (anonymous):

thus why come and i ll teach u

OpenStudy (anonymous):

ok sounds good i need the help bad

OpenStudy (zehanz):

If i is not supposed to be sqrt(-1) then everything remails the same, only the part where i²=-1 is not to be done. Then: 54+6i+36i+4i² = 4i²+42i+54 is the answer

OpenStudy (precal):